OpenStudy (jenessa_2413):

Oscar and Kim have $22.00 each to spend at a book fair, where all students receive a 30% discount. They both want to purchase a copy of the same book, which normally sells for $28.50 plus 10% sales tax. To check if he has enough to purchase the book, Oscar takes 30% of $28.50 and subtracts that amount from the normal price. He takes 10% of the discounted selling price and adds it back to find the purchase amount. Kim takes 70% of the normal purchase price and then computes 110% of the reduced price. Is Oscar correct? Is Kim correct? Do they have enough money to purchase the book? Explain your

OpenStudy (memyselfandpi):

Let's find out what each of their equations equal.

OpenStudy (memyselfandpi):

@jenessa_2413

OpenStudy (memyselfandpi):

Kim takes 70% of the normal price, which equals 15.4$, than multiplies it by 110% which equals 16.94$. That is Kim's price

OpenStudy (memyselfandpi):

Wait, I have to go... @Mehek14 can you finish this?

OpenStudy (jenessa_2413):

@Mehek14 can you help please

OpenStudy (jojo12g):

kim is right for one and yes you do have enough to buy the book. by doing it Oscars way you end up in the negative with no money to spend to get the book do it kims way you got enough and do it oscars way you dont

OpenStudy (jenessa_2413):

thanks :)

OpenStudy (jtvatsim):

I think you actually get the same answer using either Kim's or Oscar's method...

OpenStudy (jtvatsim):

For the first step, they are definitely the same: Oscar does 28.50 - 0.30*28.50 Kim does 0.70*28.50 But look! 28.50 - 0.30*28.50 = (1 - 0.30)*28.50 = 0.70*28.50 is the same algebraically

OpenStudy (jtvatsim):

They both find the discount price (without tax) to be $19.95.

OpenStudy (jtvatsim):

Now, once they go to calculate the sales tax... Oscar does 19.95 + 0.10*19.95 Kim does 1.10*19.95 But look! 19.95 + 0.10*19.95 = (1 + 0.10)*19.95 = 1.10*19.95 is the same algebraically again!

OpenStudy (jtvatsim):

The amount to be paid (using either method) is $21.95 (rounded).

OpenStudy (jenessa_2413):

so you get the same answer for both kim and oscar? im a little confused.

OpenStudy (jtvatsim):

That's what I think. If you pretend to follow their advice you will get the same answer either way.

OpenStudy (jenessa_2413):

oh ok thank you:))